OpenStudy (anonymous):

HELP!!! I will fan and give medals. I have 10 questions.

OpenStudy (anonymous):

What is the measure of each interior angle of a regular octagon? (1 point) A. 135 B. 1080 C. 147 D. 124

OpenStudy (anonymous):

Find the value of a and y. A. x= 109, y=108 B. x=108, y=109 C. x=108, y=108 D. x=109, y=109

OpenStudy (texaschic101):

Question 1 : there is a formula for this.. (180n - 360) / n --- for an octagon, n would be 8 do the math and it will tell you the measure of each angle

OpenStudy (anonymous):

So it would be c.?

OpenStudy (texaschic101):

Question 2: formula for sum of interior angles.. sum = 180(n - 2) -- n = the number of sides, there are 4. Once you find the sum, subtract all of the other angles from it to get your answer for x. for y.....71 + y = 180...solve for y first one is not C...try again, just sub into the formula I gave you

OpenStudy (anonymous):

Okay I re-worked one and got B, I used what you gave me, and for the second one I got y=109